# Minimizing False-Positive Attributions in Explanations of Non-Linear Models

### v1

@@ 3.1 Formal objective
Let $\mathbf{h}_{\mathbf{x_{\star}}}:\mathcal{X}\rightarrow\mathbb{R}^{D^{\prime}}$ be the simplified input representation used in Equation 5 and $\mathbf{w}$ be its solution. Also let $\tilde{y}=\mathbf{w}^{\!\top}\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})$ denote the local surrogate prediction. PatternLocal estimates a pattern $\mathbf{a}$ by regressing the local prediction onto the simplified input of the original data within the local neighborhood of $\mathbf{x}_{\star}$
$\mathbf{s}^{\text{PatternLocal}}(\mathbf{x_{\star}})\;=\;\mathbf{a}\;=\;\arg\!\min_{\mathbf{u}}\;\mathbb{E}_{\mathbf{x}\,\sim\,\mathcal{X}}\Bigl{[}\Pi_{\mathbf{x^{\prime}_{\star}}}(\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x}))\,\bigl{(}\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})-\mathbf{u}\,{\tilde{y}}\bigr{)}^{2}\Bigr{]}\;+\;\lambda\,Q(\mathbf{u}).$ (6)
@@
$\mathbf{a}_{\ell_{2}}\;=\;\bigl{(}\,\mathbb{E}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\big{[}\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})\,\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})^{\!\top}\bigr{]}+\lambda I\bigr{)}^{-1}\,\mathbb{E}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\big{[}\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})\,\tilde{y}\,\big{]}\;=\;\frac{\operatorname{Cov}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\!\bigl{[}\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x}),\,\tilde{y}\bigr{]}}{\operatorname{Var}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\!\bigl{[}\tilde{y}\bigr{]}\;+\;\lambda},$ (7)
where expectations, covariances, and variances are taken with respect to the normalized kernel $\Pi_{\mathbf{x^{\prime}_{\star}}}(\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x}))$ over $\mathcal{X}$. See supplementary material for derivations. Equation 7 therefore expresses $\mathbf{a}_{\ell_{2}}$ as the kernel-weighted covariance between the simplified features and the surrogate response, scaled by its regularized variance.
Figure 2: Feature–importance comparison on the XOR toy problem. We draw 2 50025002\,5002 500 i.i.d. samples from the generative process of Equation 8, so the label depends solely on the interaction between $x_{1}$ and $x_{2}$, while $x_{3}$ is a suppressor that carries no marginal predictive signal. Each plot shows $(x_{1},x_{2})$ pairs, colored by the local normalized feature importance that four XAI methods assign to the suppressor variable $x_{3}$ when applied to the smooth classifier. Every plot reports the empirical mean magnitude of this attribution across all test points. Whereas LIME, KernelSHAP, and Gradient all attribute substantial importance to $x_{3}$, PatternLocal correctly drives the attribution of $x_{3}$ to (almost) zero.
@@ 3.2 Toy example
To better illustrate the intuition behind the PatternLocal method, we present a small-scale toy example. Let $s_{1},s_{2}\sim\mathcal{N}(0,1)$ and $d\sim\mathcal{N}(0,\sigma_{d}^{2})$ be independent latent variables. We observe
$\mathbf{x}=\begin{bmatrix}x_{1}\\ x_{2}\\ x_{3}\end{bmatrix}=\begin{bmatrix}s_{1}+d\\ s_{2}-d\\ d\end{bmatrix},\qquad y=\operatorname{sign}(s_{1}s_{2}),$ (8)
@@
$\mathbf{w}=\nabla\tilde{f}(\mathbf{x}_{\star})=c_{\star}\begin{bmatrix}{x_{2}}_{\star}+{x_{3}}_{\star}\\ {x_{1}}_{\star}-{x_{3}}_{\star}\\ {x_{1}}_{\star}-{x_{2}}_{\star}-2{x_{3}}_{\star}\end{bmatrix},\qquad c_{\star}=\frac{1}{\tau}\operatorname{sech}^{2}\left(\frac{1}{\tau}\tilde{f}(\mathbf{x}_{\star})\right),$ (9)
noting that $w_{1}-w_{2}+w_{3}=0$. Therefore, we in general have $w_{3}\neq 0$, implying that local linear XAI methods (e.g., gradients, LIME, KernelSHAP) will incorrectly assign nonzero importance to the suppressor variable $x_{3}$.
@@
and hence $a_{3}=0$. Similar calculation can be done for $a_{1}$ and $a_{2}$ which yields $\mathbf{a}\propto[w_{1},w_{2},0]^{\top}$. See supplementary material for derivations. We conclude that PatternLocal correctly assigns zero feature importance to the suppressor variable $x_{3}$, in contrast to standard local linear methods. This is also confirmed with simulation as seen in Figure 2.
@@ 4 Experiments
The experimental workflow begins with the construction of synthetic image datasets whose pixel-level ground-truth attribution maps are known by design, giving us full control over class-conditional distributions [^1] Code is available at https://github.com/gjoelbye/PatternLocal. Each candidate classifier is trained on a standard train/validation split and its performance is subsequently assessed on an unseen test set. For every test image, the selected XAI method produces a normalized importance map in the range $[-1,1]$. These maps are then quantitatively compared with the ground-truth attributions using the evaluation metrics introduced below. All XAI hyperparameters are tuned on the validation split before the final assessment to ensure a fair comparison.
@@ 4.1 Datasets
We focus on the XAI-TRIS benchmark dataset [7] and the artificial lesion MRI dataset [16]. Both datasets are specifically designed to assess XAI methods in the context of suppressor variables.
XAI-TRIS Benchmark
We follow the methodology outlined in Clark et al. [7]. Full details are provided in supplementary materials. The XAI-TRIS benchmark dataset consists of $64\times 64$ images (with $N=40{,}000$ samples) generated by combining a class-dependent foreground tetromino shape with background noise. The relative weighting between signal and noise is set by $\alpha$. Two noise scenarios are considered:
@@
The set of important (ground-truth) pixels is determined by the tetromino shape. For training and evaluation, each dataset is split into $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{val}}$, and $\mathcal{D}_{\text{test}}$ in a 90/5/5 ratio. Figure 3 shows examples of all scenarios. Spatially correlated noise in the CORR setting introduces suppressor pixels, background locations correlated with the foreground through the smoothing operator $G$. Although these pixels carry no direct class information, the model can exploit them to denoise the true signal [11, 28].
(b) Scenario WHITE with $\alpha=0.2$, where noise is independent (white), and no suppressor variable is present.
Artificial lesion MRI dataset
[…]
@@ 4.2 Models
We consider two types of classifiers; a multi-layer perceptron (MLP) and a convolutional neural network (CNN). Each classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$ is trained on the training dataset $\mathcal{D}_{\text{train}}$. We train the model with early stopping based on the validation performance on $\mathcal{D}_{\text{val}}$ and then we evaluate the model of the test set $\mathcal{D}_{\text{test}}$. We consider a classifier to have generalized the given classification problem when the resulting test accuracy is above 90%. See supplementary material for details on model architecture, training, and performance.
@@ 4.3 XAI methods
We evaluate 11 methods on the models, including the PatternLocal approach and a baseline. The interpretability methods applied are: LIME [17], Integrated Gradients [24], Saliency [21], DeepLift [20], GradientShap [9], and GuidedBackProp [23]. Additionally, we incorporate two edge detection filters, Sobel and Laplace, which have been shown to outperform more advanced XAI techniques in suppressor variable scenarios [7]. Note that KernelSHAP is a special case of LIME. Since we perform extensive hyperparameter optimization, KernelSHAP may also be selected during this process.
For the XAI-TRIS dataset we benchmark LIME and PatternLocal with three input-simplification schemes for $\mathbf{h}_{\mathbf{x}_{\star}}(\cdot)$: the identity mapping, a super-pixel representation, and a low-rank approximation. The same three variants are evaluated on the artificial-lesion MRI dataset. Super-pixels are generated on a uniform grid for XAI-TRIS, whereas the slic algorithm is used for the MRI data.
@@ 4.4 Metrics
The Earth Mover’s Distance (EMD) measures the optimal cost required to transform one distribution into another. For a continuous-valued importance map $s$ and ground truth $F^{+}$, both normalized to have the same mass, the EMD is computed using the Euclidean distance as the ground metric. The optimal transport cost, $\text{OT}(|\mathbf{s}|,F^{+})$, is calculated following the algorithm proposed by Bonneel et al. [6] and implemented in the Python Optimal Transport library [10]. We define a normalized EMD performance score as
@@
$\text{IME}=1-\sum_{i=1}^{|\mathcal{F}^{+}|}|s_{i}|/\sum_{i=1}^{|\mathbf{s}|}|s_{i}|.$ (11)
@@ 4.5 Hyperparameter optimization
To ensure a fair evaluation of the XAI methods, we perform computationally intensive hyperparameter optimization for each uniquely generated dataset, scenario, and signal-to-noise ratio (SNR), as well as for the corresponding trained machine learning models. We use Bayesian optimization with the Tree of Parzen Estimators (TPE) algorithm [4], which is well-suited for handling the mix of continuous and categorical hyperparameters in our setup. This optimization is carried out using the hyperopt Python package [5]. The loss function we minimize is the Earth Mover’s Distance (EMD) between the absolute value of the normalized explanation and the ground truth. A detailed description of the hyperparameter selection process for the XAI methods is provided in the supplementary material.
5 Results
XAI-TRIS Benchmark
@@ 5 Results
Figure 4: Quantitative evaluation of feature importance maps for the MLP model using $64\times 64$ XAI-TRIS benchmark images generated by various methods, as a function of the SNR. The MLP model achieves at least $90\%$ accuracy, and method hyperparameters were optimized for EMD. The function $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$ is the identity function, and $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$. The top panel (4) shows results based on EMD, while the bottom panel (4) presents performance in terms of IME. PatternLocal demonstrates significant improvements over most methods. Shaded regions indicate standard deviation as the standard error is too small to be visible.
@@
We evaluate the performance of the XAI methods on the XAI-TRIS benchmark both qualitatively and quantitatively. Qualitative evaluation is done through visual inspection, which is standard in XAI research [e.g., 3, 15, 7], and illustrated in supplementary material. The quantitative evaluation uses ground truth along with the EMD and IME metrics, as shown in Figure 4. The results presented here are for the MLP model, $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$ as the identity function and $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$. Results for the CNN model, $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$ as superpixels or low-rank approximation, and for $R(\cdot)=Q(\cdot)=\|\cdot\|_{1}$, are provided in the supplementary material.
@@
Figure 5: Examples from the artificial lesion MRI dataset, showing two MRI slices (rows). The first column displays the original MRI slices with artificially generated lesions; the second column (Ground Truth) shows the true lesion locations. The subsequent columns show feature importance maps from PatternLocal, LIME, Sobel, and Laplace. We use a CNN model, $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$, with superpixels, and set $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$ for both PatternLocal and LIME. Note that the filter-based methods fail in these examples due to the lack of clear edges. Overlaid on the ground truth are superpixels with feature importance above 0.9 from PatternLocal (green) and LIME (red). PatternLocal explanations appear better aligned with the lesion locations. The overlap is not perfect, as it is limited by the superpixel resolution.
@@
In Figure 4, we present the aggregated results of the methods across the test set for the EMD and IME metrics. Overall, the PatternLocal method significantly outperforms all other methods in the XOR and RIGID scenarios. However, in the RIGID CORR scenario, the results are mixed: PatternLocal performs substantially better than other XAI methods but is comparable to or worse than the filter methods Sobel and Laplace. These filter methods also performed well in Clark et al. [7], which is due to the XAI-TRIS methods involving rigid bodies with sharp lines, making them especially easy for filter-based methods to detect. This advantage does not carry over to images with more complex backgrounds, such as those in the artificial lesion MRI dataset.
@@
Additional ablation studies on the impact of hyperparameters and the role of the suppressor variable are provided in the supplementary material.
Artificial lesion MRI dataset
@@
We evaluate the XAI methods using the same procedure on the artificial lesion MRI dataset. The results reported here correspond to the CNN model, with $\mathbf{h}_{\mathbf{x}_{\star}}(\cdot)$ defined over superpixels and $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$. Figure 5 illustrates the performance of PatternLocal in comparison to LIME and the filter-based methods Sobel and Laplace. Additional qualitative and quantitative benchmark results are provided in the supplementary material.

### v3

@@ 3.1 Formal objective
Let $\mathbf{h}_{\mathbf{x_{\star}}}:\mathcal{X}\rightarrow\mathbb{R}^{D^{\prime}}$ be the simplified input representation used in Eq.˜5 and $\mathbf{w}$ be its solution. Also let $\tilde{y}=\mathbf{w}^{\!\top}\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})$ denote the local surrogate prediction. PatternLocal estimates a pattern $\mathbf{a}$ by regressing the local prediction onto the simplified input of the original data within the local neighborhood of $\mathbf{x}_{\star}$
$\mathbf{s}^{\text{PatternLocal}}(\mathbf{x_{\star}})\;=\;\mathbf{a}\;=\;\arg\!\min_{\mathbf{u}}\;\mathbb{E}_{\mathbf{x}\,\sim\,\mathbb{P}_{\mathcal{X}}}\Bigl[\Pi_{\mathbf{x^{\prime}_{\star}}}(\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x}))\,\bigl\|\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})-\mathbf{u}\,{\tilde{y}}\bigr\|^{2}_{2}\Bigr]\;+\;\lambda\,Q(\mathbf{u}).$ (6)
@@
$\mathbf{a}_{\ell_{2}}\;=\;\bigl(\,\mathbb{E}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\big[\tilde{y}^{2}\bigr]+\lambda I\bigr)^{-1}\,\mathbb{E}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\big[\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})\,\tilde{y}\,\big]\;=\;\frac{\operatorname{Cov}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\!\bigl[\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x}),\,\tilde{y}\bigr]}{\operatorname{Var}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\!\bigl[\tilde{y}\bigr]\;+\;\lambda},$ (7)
where expectations, covariances, and variances are taken with respect to the normalized kernel $\Pi_{\mathbf{x^{\prime}_{\star}}}(\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x}))$ over $\mathcal{X}$. See Appendix B.1 for derivations. Eq.˜7 therefore expresses $\mathbf{a}_{\ell_{2}}$ as the kernel-weighted covariance between the simplified features and the surrogate response, scaled by its regularized variance. Also see Appendix B.3 for more on sample complexity.
Figure 2: Feature–importance comparison on the XOR toy problem. We draw 2 5002\,500 i.i.d. samples from the generative process of Eq.˜8, so the label depends solely on the interaction between $x_{1}$ and $x_{2}$, while $x_{3}$ is a suppressor that carries no marginal predictive signal. Each plot shows $(x_{1},x_{2})$ pairs, colored by the local normalized feature importance that four XAI methods assign to the suppressor variable $x_{3}$ when applied to the smooth classifier. Every plot reports the empirical mean magnitude of this attribution across all test points. Whereas LIME, KernelSHAP, and Gradient all attribute substantial importance to $x_{3}$, PatternLocal correctly drives the attribution of $x_{3}$ to (almost) zero.
@@ 3.2 Toy example
We present a small-scale toy example to better illustrate the intuition behind the PatternLocal method. Let $m_{1},m_{2}\sim\mathcal{N}(0,1)$ and $d\sim\mathcal{N}(0,\sigma_{d}^{2})$ be independent latent variables. We observe
$\mathbf{x}=\begin{bmatrix}x_{1}\\ x_{2}\\ x_{3}\end{bmatrix}=\begin{bmatrix}m_{1}+d\\ m_{2}-d\\ d\end{bmatrix},\qquad y=\operatorname{sign}(m_{1}m_{2}),$ (8)
@@
$\mathbf{w}=\nabla{f_{\tau}}(\mathbf{x}_{\star})=c_{\star}\begin{bmatrix}{x_{2}}_{\star}+{x_{3}}_{\star}\\ {x_{1}}_{\star}-{x_{3}}_{\star}\\ {x_{1}}_{\star}-{x_{2}}_{\star}-2{x_{3}}_{\star}\end{bmatrix},\qquad c_{\star}=\frac{1}{\tau}\operatorname{sech}^{2}\left(\frac{1}{\tau}\tilde{f}(\mathbf{x}_{\star})\right),$ (9)
noting that $w_{1}-w_{2}+w_{3}=0$. Therefore, we generally have $w_{3}\neq 0$, implying that local linear XAI methods (e.g., gradients, LIME, KernelSHAP) will incorrectly assign nonzero importance to the suppressor variable $x_{3}$.
@@
and hence $a_{3}=0$. Similar calculation can be done for $a_{1}$ and $a_{2}$ which yields $\mathbf{a}\propto[w_{1},w_{2},0]^{\top}$. See Appendix B.2 for derivations. We conclude that PatternLocal correctly assigns zero feature importance to the suppressor variable $x_{3}$, in contrast to standard local linear methods. This is also confirmed with simulation as seen in Figure˜2.
@@ 4 Experiments
The experimental workflow begins with the construction of synthetic image datasets whose pixel-level ground-truth attribution maps are known by design, giving us complete control over class-conditional distributions. [^1] Code is available at https://github.com/gjoelbye/PatternLocal. We train each candidate classifier on a standard train/validation split and assess its performance on an unseen test set. For every test image, the selected XAI method produces a normalized importance map in the range [1,1][-1,1]. These maps are then quantitatively compared with the ground-truth attributions using the evaluation metrics introduced below. All XAI hyperparameters are tuned on the validation split before the final assessment to ensure a fair comparison.
(b) Scenario WHITE with $\alpha=0.2$, where noise is independent (white) thus no suppressor variable.
@@ 4.1 Datasets
We focus on the XAI-TRIS benchmark dataset [8] and the artificial lesion MRI dataset [19]. Both datasets specifically enable the assessment of XAI methods in the context of suppressor variables.
XAI-TRIS Benchmark.
We follow the methodology outlined in Clark et al. [8] and provide complete details in Appendix C.1. The XAI-TRIS benchmark dataset consists of $64\times 64$ images (with N=40, [^0] N=40{,}000 samples) generated by combining a class-dependent foreground tetromino shape with background noise. The relative weighting between signal and noise is set by $\alpha$. We consider two noise scenarios:
@@
The tetromino shape determines the set of important (ground-truth) pixels. For training and evaluation, each dataset is split into $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{val}}$, and $\mathcal{D}_{\text{test}}$ in a 90/5/5 ratio. Figure 3 shows examples of all scenarios. Spatially correlated noise in the CORR setting introduces suppressor pixels, background locations correlated with the foreground through the smoothing operator $G$. Although these pixels carry no direct class information, the model can exploit them to denoise the true signal [12, 34].
Artificial lesion MRI dataset.
Following Oliveira et al. [19], the Artificial Lesion MRI dataset contains grayscale medical downscaled images, with $D=128\times 128=16\,384$. It includes N=7 500N=7\,500 samples, evenly split into $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{val}}$, and $\mathcal{D}_{\text{test}}$. Each sample simulates a brain MRI slice with artificial lesions of predefined shapes and intensities, loosely resembling real-world white matter hyperintensities (WMH). Unlike XAI-TRIS, this dataset introduces challenges with more complex, organic lesion structures and anatomical background variations, making it a valuable benchmark for neuroimaging explainability. See Appendix C.2 for more details.
EEG Motor Imagery dataset.
We use the PhysioNet EEG Motor Movement/Imagery dataset [24], containing recordings from 109 subjects performing motor tasks. EEG was acquired from 64 scalp electrodes at 160 Hz following the 10–10 system. We focus on runs R03, R07, and R11, where subjects repeatedly opened and closed their left or right fist. Labels T1 and T2 mark left- and right-fist movements, defining a binary classification problem ($y=1$ for left, $y=0$ for right). We use 19 standard 10–20 electrodes, resample to 100 Hz, apply a 0.5 Hz high-pass filter, and re-reference to the common average. A 3s window (1s before to 2s after onset) is extracted for each event, yielding 19 channels and 300 time points per trial. After removing invalid segments, the final dataset comprises N=4,927N=4{,}927 balanced trials (2,456 right, 2,471 left), split into $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{val}}$, and $\mathcal{D}_{\text{test}}$ in a 90/5/5 ratio.
@@ 4.2 Models
We use two types of classifiers for the synthetic datasets: a multi-layer perceptron (MLP) and a convolutional neural network (CNN). For the EEG Motor Imagery dataset, we use the ShallowFBCSPNet from Schirrmeister et al. [25]. Each classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$ is trained on the training dataset $\mathcal{D}_{\text{train}}$. We train the models with early stopping based on validation performance on $\mathcal{D}_{\text{val}}$, and then evaluate them on the test set $\mathcal{D}_{\text{test}}$. We consider a classifier to have generalized to the classification problem if its test accuracy exceeds 90%. See Appendix D.1 for details on model architecture, training, and performance.
@@ 4.3 XAI methods
We evaluate nine methods on the models, including the PatternLocal approach and baselines. The XAI methods applied are: LIME [21], Integrated Gradients [30], Saliency [27], DeepLift [26], GradientShap [10], and GuidedBackProp [29]. Additionally, we incorporate two edge detection filters, Sobel and Laplace, which prior work shows outperform more advanced XAI techniques in suppressor variable scenarios [8]. Note that KernelSHAP is a special case of LIME. Since we perform extensive hyperparameter optimization, we may also select KernelSHAP during this process.
For the XAI-TRIS dataset, we benchmark LIME and PatternLocal with three input-simplification schemes for $\mathbf{h}_{\mathbf{x}_{\star}}(\cdot)$: the identity mapping, a superpixel representation, and a low-rank approximation. We evaluate the same three variants on the artificial-lesion MRI dataset. Superpixels are generated on a uniform grid for XAI-TRIS, whereas we use the slic algorithm for the MRI dataset.
@@ 4.4 Metrics
We evaluate explanation quality using two complementary metrics: the Earth Mover’s Distance (EMD) and the Importance Mass Error (IME). See Appendix F.1 for a discussion of additional metrics such as faithfulness.
@@
The Earth Mover’s Distance (EMD) measures the optimal cost to transform one distribution into another. For a continuous-valued importance map $s$ and ground truth $\mathcal{F}^{+}$, both normalized to have the same mass, we compute the EMD using the Euclidean distance as the ground metric. The optimal transport cost, $\text{OT}(|\mathbf{s}|,\mathcal{F}^{+})$, is calculated following the algorithm proposed by Bonneel et al. [7] and implemented in the Python Optimal Transport library [11]. We define a normalized EMD performance score as
@@
$\text{IME}=1-\sum_{i=1}^{\operatorname{card}(\mathcal{F}^{+})}|s_{i}|/\sum_{i=1}^{\operatorname{card}(\mathbf{s})}|s_{i}|.$ (11)
@@ 4.5 Hyperparameter optimization
For both the synthetic XAI-TRIS Benchmark and the semi-synthetic Artificial Lesion MRI dataset, we conduct extensive hyperparameter optimization across all datasets, scenarios, SNR levels, and corresponding models to ensure a fair evaluation of all XAI methods. We employ Bayesian optimization using the Tree-structured Parzen Estimator (TPE) algorithm [4], implemented via the hyperopt package [5], which efficiently manages mixed continuous and categorical search spaces. The optimization objective minimizes the Earth Mover’s Distance (EMD) between the absolute normalized explanation and the ground truth on the validation set. Additional implementation details are provided in Appendix D.2.
@@
Figure 4: Quantitative evaluation of feature importance maps for the MLP model using $64\times 64$ XAI-TRIS benchmark images generated by various methods, as a function of the SNR. The MLP model achieves at least $90\%$ accuracy, and the method’s hyperparameters were optimized for EMD. The function $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$ is the identity function, and $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$. The top panel (4) shows results based on EMD, while the bottom panel (4) presents performance in terms of IME. PatternLocal demonstrates significant improvements over most methods. Shaded regions indicate standard deviation.
5 Results
XAI-TRIS Benchmark.
@@ 5 Results
We evaluate the performance of the XAI methods on the XAI-TRIS benchmark both qualitatively and quantitatively. We conduct qualitative evaluation through visual inspection, which is standard in XAI research [e.g., 3, 17, 8], and illustrate it in Appendix E.1. The quantitative evaluation uses ground truth along with the EMD and IME metrics, as shown in Figure 4. The results presented here are for the MLP model, $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$ as the identity function and $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$. Results for the CNN model, $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$ as superpixels or low-rank approximation, and for $R(\cdot)=Q(\cdot)=\|\cdot\|_{1}$, are provided in Appendix E.1.
@@
Figure 5: Examples from the artificial lesion MRI dataset, showing two MRI slices (rows). The first column displays the original MRI slices with artificially generated lesions; the second column (Ground Truth) shows the true lesion locations. The subsequent columns show feature importance maps from PatternLocal, LIME, Sobel, and Laplace. We use a CNN model, $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$, with superpixels, and set $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$ for both PatternLocal and LIME. Note that the filter-based methods fail in these examples due to the lack of clear edges. Overlaid on the ground truth are superpixels with feature importance above 0.9 from PatternLocal (green) and LIME (red). PatternLocal explanations appear better aligned with the lesion locations. The overlap is not perfect, as the superpixel resolution limits it.
Figure 6: (a) Channel-averaged explanation signals of the PatternLocal explanation shows a marked increase around cue onset. (b-c) First SSD patterns derived from the explanation in the alpha (8-12 Hz) and gamma (30-50 Hz) bands have a good dipole fit (alpha r = 0.92, gamma r = 0.84). (d-e) eLORETA source localization of the SSD-reconstructed explanation highlights peri-rolandic cortex with stronger power in the left hemisphere, consistent with contralateral activation during right-hand imagery. SSD is applied directly to single-trial explanations before source reconstruction.
@@
In Figure 4, we present the aggregated results of the methods across the test set for the EMD and IME metrics. Overall, the PatternLocal method substantially outperforms all other methods in the XOR and RIGID scenarios. However, we observe mixed results in the RIGID CORR scenario: PatternLocal performs substantially better than other XAI methods but is comparable to or worse than the filter methods Sobel and Laplace. These filter methods also performed well in Clark et al. [8] due to XAI-TRIS images having rigid bodies with sharp lines, making them especially easy for filter-based methods to detect. This advantage does not apply to images with more complex backgrounds, such as those in the artificial lesion MRI dataset.
@@
Additional ablation studies on hyperparameters’ impact and the suppressor variable’s role are in Appendix G.
Artificial lesion MRI dataset.
@@
We evaluate the XAI methods using the same procedure on the artificial lesion MRI dataset. The results reported here correspond to the CNN model, with $\mathbf{h}_{\mathbf{x}_{\star}}(\cdot)$ defined over superpixels and $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$. Figure˜5 illustrates the performance of PatternLocal compared to LIME and the filter-based methods Sobel and Laplace. We provide additional qualitative and quantitative benchmark results in Appendix E.2.
EEG Motor Imagery dataset.
Since the EEG motor imagery dataset does not have a ground-truth, we assess the explanations’ physiological plausibility. This didactic example demonstrates how explanations can be meaningfully analyzed when they align with the underlying neurophysiology.
We used PatternLocal with the bandwidth set to the median pairwise distance, $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$ as the identity function and no regularization. Applying PatternLocal to the motor-imagery EEG data produced explanations with clear temporal and spatial structure. The explanations concentrate on task onset and retain distinct oscillatory patterns in the alpha and gamma bands. When we apply spatio-spectral decomposition (SSD) [18] to single-trial explanations, the resulting components remain physiologically interpretable and localize with eLORETA[20] to the expected motor regions. For the right-hand example shown in Figure 6, the source reconstruction reveals increased left-hemisphere activity, reflecting contralateral motor engagement. Across the test set, dipole fits of SSD patterns indicate higher physiological plausibility for PatternLocal ($0.756\pm 0.090$), with the raw instances showing similar results ($0.738\pm 0.013$) and LIME performing worse ($0.604\pm 0.013$). These findings demonstrate that PatternLocal yields explanations consistent with the physical forward model, enabling valid time-frequency and source analyses, unlike methods such as LIME, which do not permit meaningful decomposition of their outputs in this way.